Differences between REVTEX 4 and REVTEX 3

American Physical Society^{†}

1 Research Road, Ridge, NY 11961

(Dated: May 2000)

Contents

1 Introduction 1

2 Version of L^{A}TEX 1

3 Class Options and Defaults 1

4 One- and Two-column formatting 1

5 Frontmatter Markup 1 5.1 Authors, Affiliations, and Author Notes . 2 5.2 Abstracts . . . . . . . . . . . . . . . . . . 2

6 Citations and References 2

7 Footnotes and Tablenotes 2

8 Section Commands 3

9 Figures 3

10Tables 3

11Font selection 3

12Math and Symbols 4

13Obsolete REVTEX 3.1 commands 4

14Converting a REVTEX 3.1 Document to REVTEX 4 4

1. Introduction

This document gives a brief summary of how REVTEX 4 is different from what authors may already be familiar with. The two primary design goals for REVTEX 4 are to 1) move to L^{A}TEX 2ε and 2) improve the markup so that infomation can be more reliably extracted for the editorial and production processes. Both

†Electronic address: revtex4@aps.org

of these goals require that authors comfortable with earlier versions of REVTEX change their habits. In addition, authors may already be familiar with the standard article.cls in L^{A}TEX 2ε. REVTEX 4 differs in some important ways from this class as well. For more complete documentation on REVTEX 4, see the main REVTEX 4 Author’s Guide. The most important changes are in the markup of the frontmatter (title, authors, affiliations, abstract, etc.) . Please see Sec. 5.

2. Version of L^{A}TEX

The most obvious difference between REVTEX 4 and REVTEX 3 is that REVTEX 4 works solely with L^{A}TEX 2ε; it is not useable as a L^{A}TEX2.09 package. Furthermore, REVTEX 4 requires an up-to-date L^{A}TEX installation (1996/06/01 or later); its use under older versions is not supported.

3. Class Options and Defaults

Many of the class options in REVTEX 3 have been retained in REVTEX 4. However, the default behavior for these options can be different than in REVTEX 3. Currently, there is only one society option, aps, and this is the default. Furthermore, the selection of a journal (such as prl) will automatically set the society as well (this will be true even after other societies are added).

In REVTEX 3, it was necessary to invoke the floats, but this is automatic in REVTEX 4. REVTEX 4 introduces two new class options, endfloats and endfloats* for moving floats to the end of the paper.

The preamble commands \draft and \tighten have been replaced with new class options draft and tightenlines, respectively. The \preprint command is now used only for specifying institutional report numbers (typeset in the upper-righthand corner of the first page); it no longer influences whether PACS numbers are displayed below the abstract. PACS display is controlled by the showpacs and noshowpacs (default) class options.

Paper size options (letter, a4paper, etc.) work in REVTEX 4. The text “Typeset by REVTEX” no longer appears by default - the option byrevtex will place this text in the lower-lefthand corner of the first page.

4. One- and Two-column formatting

REVTEX 4 has excellent support achieving the twocolumn formatting in the Phys. Rev. style. It will balance the columns automatically. Whereas REVTEX 3 had the \widetext and \narrowtext commands for switching between one- and two-cloumn modes, REVTEX 4

simply has a widetext environment, \begin{widetext}

… \end{widetext}. One-column formatting can be specified by choosing either the onecolumn or preprint class option (the REVTEX 3 option manuscript no longer exists). Two-column formatting is the default for most journal styles, but can be specified with the twocolumn option. Note that the spacing for preprint is now set to 1.5, rather than full double-spacing. The tightenlines option can be used to reduce this to single spacing.

5. Frontmatter Markup

REVTEX 4 has substantially changed how the frontmatter for an article is marked up. These are the most significant differences between REVTEX 4 and other systems for typesetting manuscripts. It is essential that authors new to REVTEX 4 be familiar with these changes.

5.1 Authors, Affiliations, and Author Notes

REVTEX 4 has substantially changed the markup of author names, affiliations, and author notes (footnotes giving additional information about the author such as a permanent address or an email address).

• Each author name should appear separately in individual \author macros.

• Email addresses should be marked up using the \email macro.

• Alternative affiliation information should be marked up using the \altaffiliation macro.

• URLs for author home pages can be specified with a

  \homepage macro.

• The \thanks macro should only be used if one of the above don’t apply.

• The grouping of authors by affiliations is accomplished automatically. Each affiliation should be in its own

  \affiliation command. Multiple \affiliation,
  \email, \homepage, \altaffiliation, and \thanks

commands can be applied to each author. The macro \and has been eliminated.

• Footnote-style associations of authors with affilitations should not be done via explicit superscripts; rather, the class option superscriptaddress should be used to accomplish this automatically.

• A collaboration for a group of authors can be given using the \collaboration command.

• \affiliation commmands apply to all previous authors that don’t have an affiliation already declared. Furthermore, for any particular author, the

                                                    \affilation must follow any \email, \homepage,
                                                    \altaffiliation, or \thanks commands for that au-

thor.

                                                  • \email, \homepage, \altaffiliation, and \thanks

commands are grouped together under a single footnote for each author. These footnotes can either appear at the bottom of the first page of the article or as the first entries in the bibliography. The journal style controls this placement, but it may be overridden by using the class options bibnotes and nobibnotes. Note that these footnotes are treated differently than the other footnotes in the article.

Table I summarizes some common mistakes.

5.2 Abstracts

REVTEX 4, like REVTEX 3, uses the abstract envi-

                                                  ronment \begin{abstract} . . . \end{abstract} for the

abstract. The abstract environment must appear before the \maketitle command in REVTEX 4. The abstract will be formatted appropriately for either one-column (preprint) or two-column formatting. In particular, in the two-column case, the abstract will automatically be placed in a single column that spans the width of the page. It is unnecessary to use a \minipage or any other macro to achieve this result.

6. Citations and References

REVTEX 4 uses the same \cite,\ref, and \bibitem commmands as standard L^{A}TEX and REVTEX 3. Citation handling is based upon Patick Daly’s natbib package. The references environment is no longer used. Instead, use the standard L^{A}TEX 2ε environment

                                                  thebibliography.

Two new BibTEX files have been included with REVTEX 4, apsrev.bst and apsrmp.bst. These will format references in the style of Physical Review and Reviews of Modern Physics respectively. In addition, these BibTEX styles automatically apply a special macro \bibinfo to each element of the bibliography to make it easier to extract information for use in the editorial and production processes. Authors are strongly

REVTEX 3 Markup  REVTEX 4 Markup  Explanation

\author{Author One and Author Two}            \author{Author One}  One name per
\author{Author Two}                 \author
\author{Author One$^{1}$}                     \author{Author One}  Use superscriptaddress

…  …  class option

\affiliation{$^{1}$APS}                       \affiliation{APS}
\thanks{Permanent address...}                 \altaffiliation{}  Use most specific macro
\thanks{Electronic address: user@domain.edu}  \email{user@domain.edu}  available
\thanks{http://publish.aps.org/}              \homepage{http://publish.aps.org/}

TABLE I: Common mistakes in marking up the frontmatter

urged to use BibTEX to manage their bibliographies so that the \bibinfo directives will be automatically included. Other bibliography styles can be specified by using the \bibliographystyle command, but unlike standard L^{A}TEX 2ε, you must give this command before the

\begin{document} statement.

Please note that the package cite.sty is not needed with REVTEX 4 and is incompatible.

7. Footnotes and Tablenotes

REVTEX 4 uses the standard \footnote macro for footnotes. Footnotes can either appear on the bottom of the page on which they occur or they can appear as entries at the end of the bibliography. As with author notes, the journal style option controls the placement; however, this can be overridden with the class options

footinbib and nofootinbib.

Within a table, the \footnote command behaves differently. Footnotes appear at the bottom of the table.

\footnotemark and \footnotetext are also available

within the table environment so that multiple table entries can share the same footnote text. There is no

longer a need to use a \tablenote, \tablenotemark,
and \tablenotetext macros.

8. Section Commands

The title in a \section{⟨title⟩} command will be automatically uppercased in REVTEX 4. To prevent a particular letter from being uppercased, enclose it in curly braces.

9. Figures

Figures should be enclosed within either a figure or figure* environment (the latter will cause the figure to span the full width of the page in two-column mode). L^{A}TEX 2ε has two convenient packages for including the figure file itself: graphics and graphicx. These two packages both define a macro \includegraphics which calls in the figure. They differ in how arguments for rotation, translation, and scaling are specified. The package epsfig has been re-implemented to use these graphicx package. The package epsfig provides an interface similar to that under the REVTEX 3 epsf class option. Authors should use these standard L^{A}TEX 2ε packages rather than some other alternative.

10. Tables

Tables should be enclosed within either a table or table* environmnent (the latter will cause the table to span the full width of the page in two-column mode). The heart of the table is the tabular environment. This will behave for the most part as in standard L^{A}TEX 2ε. Note that REVTEX 4 no longer automatically adds double (Scotch) rules around tables. Nor does the tabular environment set various table parameters as before. Instead, a new environment ruledtabular provides this functionality. This environment should surround the tabular environment:

\begin{table}
\caption{...}
\label{tab:...}
\begin{ruledtabular}
\begin{tabular}
...
\end{tabular}
\end{ruledtabular}
\end{table}

Under REVTEX 3, tables automatically break across pages. REVTEX 4 provides some of this functionality. However, this requires adding the table a float placement option of [h] or [H] (meaning put the table “here”) to the \begin{table} command. Furthermore, if the ruledtabular environment is not used, the \begin{tabular} command must be given an optional [v] argument. Tables set in the table* environment ignore any float placement options.

Long tables are more robustly handled by using the longtable.sty package included with the standard L^{A}TEX 2ε distribution (put \usepackage{longtable} in the preamble). This package gives precise control over the layout of the table. REVTEX 4 goes out of its way to provide patches so that the longtable environment will work within a two-column format. A longtable that is too wide for the narrow columns can always be set within a widetext environment. (Note that a standard table should never be enclosed in a widetext environment - use the table* enviroment instead.)

To create tables with columns of numbers aligned on decimal points, load the standard L^{A}TEX 2ε dcolumn package and use the d column specifier. The content of each cell in the column is implicitly in math mode: Use of math delimiters ($) is unnecessary in a d column.

Footnotes within a table can be specified with the \footnote command (see Sec. 7).

11. Font selection

The largest difference between REVTEX 3 and REVTEX 4 with respect to fonts is that REVTEX 4 allows one use the L^{A}TEX 2ε font commands such as \textit, \texttt, \textbf etc. This commands should be used in place of the basic TEXL^{A}TEX 2.09 font commands such as \it, \tt, \bf, etc. The new font commands better handle subtleties such as italic correction and scaling in super- and subscripts.

12. Math and Symbols

REVTEX 4 depends more heavily on packages from the standard L^{A}TEX 2ε distribution and AMSL^{A}TEX than REVTEX 3 did. Thus, REVTEX 4 users should make sure their L^{A}TEX 2ε distributions are up to date and they should install AMSL^{A}TEX as well. In general, if any fine control of equation layout, special math symbols, or other specialized math constructs are needed, users should look to the amsmath package (see the AMSL^{A}TEX documentation).

REVTEX 4 provides a small number of additional diacritics, symbols, and bold parentheses. Table II summarizes this.

Here is a partial list of the more notable changes between REVTEX 3 and REVTEX 4 math:

• Bold math characters should now be handle via the standard L^{A}TEX 2ε bm package (use \bm instead of \bbox). \bm will handle Greek letters and other symbols.

• Use the class options amsmath, amsfonts and amssymb to get even more math fonts and symbols. \mathfrak TABLE II: Special REVTEX 4 symbols, accents, and boldfaced parentheses defined in revsymb.sty

                                                    \lambdabar         λ  \openone  11
                                                          \altsuccsim  ≻_{∼}  \altprecsim  ≺_{∼}
                                                          \alt  ≲  \agt  ≳
                                                          \tensor x  x^{↔}  \overstar x  x^{∗}
                                                          \loarrow x  x^{←}  \roarrow x  x^{→}
                                                          \biglb ( \bigrb)  ( ( () ) ) \Biglb ( \Bigrb)  ( ( () ) )

                                                          \bigglb ( \biggrb) ( ( () ) ) \Bigglb ( \Biggrb ) ^{( ( (}^{) ) )}

and \mathbb will, for instance, give Fraktur and Blackboard Bold symbols.

• Use the fleqn class option for making equation flush left or right. \FL and \FR are no longer provided.

• In place of \eqnum, load the amsmath package

                                                    [\usepackage{amsmath}] and use \tag.
                                                  • In place of \case, use \textstyle\frac.

• In place of the mathletters environment, load the amsmath package and use subequations environment.

                                                  • In place of \slantfrac, use \frac.
                                                  • The marcros \corresponds, \overdots, and

\overcirc have been removed. See Table III.

13. Obsolete REVTEX 3.1 commands

Table III summarizes more differences between REVTEX 4 and REVTEX 3, particularly which REVTEX 3 commands are now obsolete.

14. Converting a REVTEX 3.1 Document to REVTEX 4

REVTEX 3 documents can be converted to REVTEX 4 rather straightforwardly. The following checklist covers most of the major steps involved.

                                                  • Change  \documentstyle{revtex}  to
                                                    \documentclass{revtex4}, and run the document

under L^{A}TEX 2ε instead of L^{A}TEX2.09.

• Replace the \draft command with the draft class option.

• Replace the \tighten command with the

                                                    tightenlines class option.

• For each \author command, split the multiple authors into individual \author commands. Remove any instances of \and.

• For superscript-style associations between authors and affiliations, remove explicit superscripts and use the

superscriptaddress class option.
• Use \affiliation instead of \address.

• Put \maketitle after the abstract environment and any \pacs commands.

• If double-ruled table borders are desired, enclose tabular enviroments in ruledtabular environments.

• Convert long tables to longtable, and load the longtable package. Alternatively, give the table an [h] float placement parameter so that the table will break automatically. Additionally, the ruledtabular or adding the optional [v] argument to the \begin{tabular} command is necessary.

• Replace any instances of the \widetext and \narrowtext commands with the widetext environment. Usually, the \begin{widetext} statement will replace the \widetext command, and the \end{widetext} statement replaces the matching

\narrowtext command.

Note in this connection that due to a curious feature of L^{A}TEX itself, REVTEX 4 having a widetext environment means that it also has a definition for the \widetext command, even though the latter cammand is not intended to be used in your document. Therefore, it is particularly important to remove all \widetext commands when converting to REVTEX 4.

• Remove all obsolete commands: \FL, \FR,

\narrowtext, and \mediumtext (see Table III).

• Replace \case with \frac. If a fraction needs to be set in text style despite being in a display equation, use the construction \textstyle\frac. Note that \frac does not support the syntax \case1/2.

                                                 • Replace \slantfrac with \frac.
                                                 • Change \frak to \mathfrak{⟨char⟩} and \Bbb to

\mathbb{⟨char⟩}, and invoke one of the class options

                                                  amsfonts or amssymb.

• Replace environment mathletters with environment

                                                  subequations and load the amsmath package.

• Replace \eqnum with \tag and load the amsmath package.

• Replace \bbox with \bm and load the bm package.

• If using the \text command, load the amsmath package.

• If using the d column specifier in tabular environments, load the dcolumn package. Under dcolumn, the content of each d column cell is implicitly in math mode: remove any $ math delimiters appearing in cells in a d column.

                                                 • Replace  \tablenote  with  \footnote,
                                                  \tablenotemark with \footnotemark,  and
                                                  \tablenotetext with \footnotetext.
                                                 • Replace  \begin{references}  with
                                                  \begin{thebibliography}{}; \end{references}
                                                  with \end{thebibliography}.

TABLE III: Differences between REVTEX 3.1 and REVTEX 4 markup

REVTEX 3.1 command  REVTEX 4 replacement

\documentstyle[⟨options⟩]{revtex}  \documentclass[⟨options⟩]{revtex4}
option manuscript                  preprint

\tighten preamble command  tightenlines class option
\draft preamble command  draft class option
\author                           \author{⟨name⟩} may appear multiple times; each signifies a new author name.

\collaboration{⟨name⟩}: Collaboration name (should appear after last \author) \homepage{⟨URL⟩}: URL for preceding author \email{⟨email⟩}: email address for preceding author \altaffiliation: alternate affiliation for preceding \author

\thanks \thanks, but use only for information not covered by \email, \homepage, or

                                  \altaffilitiation

\and  obsolete, remove this command
\address                          \affiliation{⟨institution⟩} gives the affiliation for the group of authors above

\affiliation[⟨note⟩] lets you specify a footnote to this institution \noaffiliation signifies that the above authors have no affiliation

\preprint                         \preprint{⟨number⟩} can appear multiple times, and must precede \maketitle

\pacs                             \pacs must precede \maketitle

abstract environment  abstract environment must precede \maketitle
\maketitle                        \maketitle must follow all frontmatter data commands
\narrowtext  obsolete, remove this command
\mediumtext  obsolete, remove this command
\widetext  obsolete, replace with widetext environment
\FL  obsolete, remove this command
\FR  obsolete, remove this command

\eqnum  replace with \tag, load amsmath
mathletters  replace with subequations, load amsmath

tabular environment No longer puts in doubled-rules. Enclose tabular in ruledtabular to get old

behavior.

quasitable environment  obsolete, tabular environment no longer puts in rules

references environment  replace with thebibliography{}
\case  replace with \textstyle\frac
\slantfrac  replace with \frac
\tablenote  replace with \footnote
\tablenotemark  replace with \footnotemark
\tablenotetext  replace with \footnotetext
\overcirc  Use standard L^{A}TEX 2ε \mathring
\overdots  Use \dddot with amsmath
\corresponds  Use \triangleq with amssymb
epsf class option  \usepackage{epsfig}